# On hyperbolic links associated to Eulerian cycles on ideal right-angled hyperbolic $3$-polytopes

### v2

@@ Abstract.
We consider Eulerian cycles without transversal selfintersections in $4$-valent planar graphs. We prove that any cycle of this type in the graph of an ideal right-angled hyperbolic $3$-polytope corresponds to a hyperbolic link such that its complement consists of $4$-copies of this polytope glued according to its checkerboard coloring. Moreover, this link consists of trivially embedded circles bijectively corresponding to the vertices of the polytope. For such cycles we prove that the $3$-antiprism $A(3)$ (octahedron) has exactly $2$ combinatorially different cycles, the $4$-antiprism $A(4)$ has exactly $7$ combinatorially different cycles, and these cycles correspond to $7$ cycles (perhaps combinatorially equivalent) on any polytope different from antiprisms, and any antiprism $A(k)$ has at least $2$ combinatorially different cycles. The $2$-fold branched covering space corresponding to our link is a small cover over some simple $3$-polytope. This small cover is defined by a Hamiltonian cycle on it. We show that any Hamiltonian cycle on a compact right-angled hyperbolic $3$-polytope arises in this way, while for a Hamiltonian cycle on a right-angled hyperbolic $3$-polytope of finite volume the necessary and sufficient condition is that at each ideal vertex it does not go straight. We introduce a transformation of a Eulerian cycle along conjugated vertices allowing to build new cycles from a given one.
@@
convex polytope, link, chain, ideal right-angled hyperbolic polytope, small cover
@@ 2. Basic facts
Let $G$ be a planar. Its medial graph is a new planar graph $M(G)$ with vertices bijectively corresponding to edges of $G$. Its edges arise when we walk around the boundary cycle of each face. Each vertex of this cycle corresponds to an edge of $M(G)$ connecting the vertices corresponding to successive edges of the cycle. A medial graph is $4$-valent. Is is known that a graph $G$ is a graph $G(P)$ of an ideal hyperbolic right-angled $3$-polytope $P$ if and only if it is a medial graph of some simple polytope $\widehat{P}$. Moreover, $\widehat{P}$ is defined uniquely up to passing to the dual polytope $\widehat{P}^{*}$ (see more details in [E19]).
@@
Construction 2.3 (A manifold from a checkerboard coloring).
@@
Construction 2.5 (A small cover and a link from a Hamiltonian cycle).
@@
If the components of $\partial Q\setminus\Gamma$ correspond to the pairs of colors $(1,2)$ and $(3,4)$, then $\tau=e_{1}+e_{2}=e_{3}+(e_{1}+e_{2}+e_{3})$ is the hyperelliptic involution on $N(Q,\widetilde{\Lambda}_{\Gamma})$. Moreover, the mapping $N(P,\widetilde{\Lambda}_{\Gamma})\to S^{3}$ is a $2$-fold branched covering with the following branch set (see details in [EE25, Section 4.5]). The edges of $Q$ not lying in $\Gamma$ form a perfect matching of the graph $G(Q)$ – a disjoint set of edges covering all the vertices. Then the preimage of this set in $N(Q,\widetilde{\Lambda}_{\Gamma})$ and in $S^{3}$ is a disjoint set of circles $C_{\Gamma}$, each circle glued of two copies of the corresponding edge. This link is the brach set of the covering. The detailed description of this link is given in [G24]. In [VM99S2] this construction was generalized from Hamiltonian cycles to Hamiltonian theta-graphs and Hamiltonian $K_{4}$-subgraphs.
@@ 3. Main construction
Let $P$ be an ideal right-angled hyperbolic $3$-polytope. Each nonselfcrossing Eulerian cycle $\gamma$ corresponds to a link $C_{\gamma}\subset S^{3}$ consisting of $\{\text{vertices of }P\}$ circles such that its complement $S^{3}\setminus C_{\gamma}$ is homeomorphic to the hyperbolic manifold $N(P)$ glued of $4$ copies of $P$. Moreover, $C_{\gamma}$ is the branch set of the $2$-fold brach covering $N(Q(P,\gamma),\widetilde{\Lambda}_{\Gamma_{\gamma}})\to S^{3}$.
@@
Example 3.5 can be generalized as follows. It is known that any antiprism $A(k)$ ($A(3)$ is the octahedron) is an ideal right-angled $3$-polytope (see [V17]). In Fig 3 we show two different nonselfcrossing Eulerian cycles on this polytope. The hyperbolic structure on the complement to the link corresponding to the left cycle is exactly the structure defined by W.P.Thurston in [T02, Example 6.8.7]. We learn this example due to the lectures by A.Yu.Vesnin and the diploma works by D.V. Chepakova [C23] and D.A. Tsygankov [T25]. The case of $A(3)$ is also mentioned in [V17, Section 5.1].
@@ 4. Edge-twists and the existence of nonselfcrossing Eulerian cycles
This follows from Examples 3.5, 3.6, and 3.7. ∎
@@ 5. Transformations of nonselfcrossing Eulerian cycles
Construction 5.4 (Transformation of a nonselfcrossing Eulerian cycle along conjugated vertices).
@@ 6. Links associated to Hamiltonian cycles on right-angled $3$-polytopes
A fullerene is a simple $3$-polytope with only pentagonal and hexagonal faces. It is known that any fullerene is a right-angled hyperbolic polytope and the dodecahedron is the fullerene with minimal number of facets (see more details in [E19]). It was shown by F. Kardoš [K20] that any fullerene has a Hamiltonian cycle. Each Hamiltonian cycle on a fullerene corresponds to a hyperbolic link with hyperbolic $2$-fold brach covering space.
@@
Now consider right-angled hyperbolic $3$-polytopes of finite volume. Each finite vertex of such a polytope has valency $3$, and ideal vertices have valency $4$. Using Andreev’s theorem it can be shown that cutting off ideal vertices defines a bijection between combinatorial types of right-angled hyperbolic $3$-polytopes of finite volume and almost Pogorelov polytopes different from the $4$-prism (cube) and the $5$-prism (see [DO01, Theorem 10.3.1] and [E19, Theorem 6.5]). Moreover, all quadrangles of the resulting polytope arise from ideal vertices. A simple $3$-polytope is called an almost Pogorelov polytope, if it is different from the simplex, has no $3$-belts, and any its $4$-belt surrounds a quadrangular facet.
@@
Let $\Gamma$ be a Hamiltonian cycle on an almost Pogorelov $3$-polytope $Q$. Then shrinking to points the edges of the perfect matching complementary to $\Gamma$ gives an ideal right-angled polytope $P$ if and only if each quadrangle of $Q$ has three edges in $\Gamma$. In this case the polytope $Q$ has an induced nonselfcrossing Eulerian cycle $\gamma$ such that $\Gamma=\Gamma_{\gamma}$. For the link $C_{\gamma}$ the complement $S^{3}\setminus C_{\gamma}$ is hyperbolic and the $2$-fold branched covering space becomes hyperbolic after cutting along incompressible Klein bottles corresponding to quadrangles of $Q$.
@@
For each quadrangle of $Q$ the Hamiltonian cycle contains either two its opposite edges, or three of its edges. In the first case after shrinking the quadrangle becomes a bigon and we obtain no polytope. If each quadrangle intersects $\Gamma$ by three edges, then it intersects at vertices two edges of the complementary perfect matching, and by  [E19, Corollary 12.31] cutting off all the edges of the matching gives an almost Pogorelov polytope different from the $4$- and the $5$-prism and produces all its quadrangles. Then shrinking quadrangles to points we obtain an ideal right-angled hyperbolic $3$-polytope $P$. The same polytope we obtain by shrinking all the edges of the matching. By [E22a, Theorem 4.12] quadrangles of $Q$ correspond to incompressible Klein bottles in $N(Q,\widetilde{\Lambda}_{\Gamma})$ such that the complement to their union has a complete hyperbolic structure of finite volume. ∎
@@
To characterise ideal right-angled $3$-polytopes corresponding to Hamiltonian cycles on (a) compact right-angled hyperblolic $3$-polytopes (b) right-angled hyperblolic $3$-polytopes of finite volume.
@@
In [E19, Theorem 9.17] is was proved that any ideal right-angled hyperbolic $3$-polytope $P$ can be obtained from an almost Pogorelov polytope or the polytope $P_{8}$ (drawn in the center at the bottom of Fig. 2) by a contraction of edges of a perfect matching such that no quadrangle contains two edges of the matching. Nevertheless, the complement to this matching may be not a Hamiltonian cycle, but a union of cycles containing all the vertices of the polytope.

### v3

# On hyperbolic links associated to Eulerian subgraphs on right-angled hyperbolic $3$-polytopes of finite volume

@@ Abstract.
We consider Eulerian cycles without transversal selfintersections in $4$-valent planar graphs. We prove that any cycle of this type in the graph of an ideal right-angled hyperbolic $3$-polytope corresponds to a hyperbolic link such that its complement consists of $4$-copies of this polytope glued according to its checkerboard coloring. Moreover, this link consists of trivially embedded circles bijectively corresponding to the vertices of the polytope. For such cycles we prove that the $3$-antiprism $A(3)$ (octahedron) has exactly $2$ combinatorially different cycles, the $4$-antiprism $A(4)$ has exactly $7$ combinatorially different cycles, and these cycles correspond to $7$ cycles (perhaps combinatorially equivalent) on any polytope different from antiprisms, and any antiprism $A(k)$ has at least $2$ combinatorially different cycles. The $2$-fold branched covering space corresponding to our link is a small cover over some simple $3$-polytope. This small cover is build from a Hamiltonian cycle on this polytope by the A.D. Mednykh’s construction. We show that any Hamiltonian cycle on a compact right-angled hyperbolic $3$-polytope arises in this way, while for a Hamiltonian cycle on a right-angled hyperbolic $3$-polytope of finite volume the necessary and sufficient condition is that at each ideal vertex it does not go straight. We introduce a transformation of a Eulerian cycle along conjugated vertices allowing to build new cycles from a given one. The link corresponding to a Hamiltonian cycle on a simple $3$-polytope always contains the Hopf link consisting of two circles. We consider links corresponding to Hamiltonian theta-graphs and Hamiltonian $K_{4}$-graphs on simple $3$-polytopes introduced by A.D. Mednykh and A.Yu. Vesnin. We give a criterion when such a link consists of mutually unlinked circles and when it is trivial. We give a necessary condition for such a link to be hyperbolic. The simplest example is the Borromean rings corresponding to the Hamiltonian theta-graph on the cube. We introduce the notions of a nonselfcrossing Eulerian theta-graph and $K_{4}$-graph on a right-angled hyperbolic $3$-polytope of finite volume with $2$ or $4$ finite vertices and construct the corresponding hyperbolic link.
@@
Convex polytope, Eulerian cycle, ideal right-angled hyperbolic polytope, small cover, Borromean rings, Brunnian link
@@ 1. Introduction
The link corresponding to a Hamiltonian cycle on a simple $3$-polytope always contains the Hopf link consisting of two circles. In Section 7 we consider links corresponding to Hamiltonian theta-graphs and Hamiltonian $K_{4}$-graphs on simple $3$-polytopes introduced by A.D. Mednykh and A.Yu. Vesnin in [VM99S2]. We give a criterion when such a link consists of mutually unlinked circles and when it is trivial. We give a necessary condition for such a link to be hyperbolic. The simplest example is the Borromean rings corresponding to the Hamiltonian theta-graph on the cube.
In Section 8 we introduce the notions of a nonselfcrossing Eulerian theta-graph and $K_{4}$-graph on a right-angled hyperbolic $3$-polytope of finite volume with $2$ or $4$ finite vertices and construct the corresponding hyperbolic link.
@@ 2. Basic facts
Remark 2.2.
In the paper [BFFS18] devoted to Barnette’s conjecture that every $3$-connected cubic planar bipartite graph is Hamiltonian, a nonselfcrossing Eulerian cycle is called an $A$-trail.
Remark 2.3.
As it was mentioned by D.V. Talalaev in private communication nonselfcrossing Eulerian cycles in $4$-valent planar graphs play an important role in the theory of electrical networks [BKT26].
@@
Let $G$ be a planar. Its medial graph is a new planar graph $M(G)$ with vertices bijectively corresponding to edges of $G$. Its edges arise when we walk around the boundary cycle of each face. Each vertex of this cycle corresponds to an edge of $M(G)$ connecting the vertices corresponding to successive edges of the cycle. A medial graph is $4$-valent. It is known that a graph $G$ is a graph $G(P)$ of an ideal hyperbolic right-angled $3$-polytope $P$ if and only if it is a medial graph of some simple polytope $\widehat{P}$. Moreover, $\widehat{P}$ is defined uniquely up to passing to the dual polytope $\widehat{P}^{*}$ (see more details in [E19]).
@@
Construction 2.5 (A manifold from a checkerboard coloring).
@@
Construction 2.7 (A small cover and a link from a Hamiltonian cycle).
@@
If the components of $\partial Q\setminus\Gamma$ correspond to the pairs of colors $(1,2)$ and $(3,4)$, then $\tau=e_{1}+e_{2}=e_{3}+(e_{1}+e_{2}+e_{3})$ is the hyperelliptic involution on $N(Q,\widetilde{\Lambda}_{\Gamma})$. Moreover, the mapping $N(P,\widetilde{\Lambda}_{\Gamma})\to S^{3}$ is a $2$-fold branched covering with the following branch set (see details in [EE25, Section 4.5]). The edges of $Q$ not lying in $\Gamma$ form a matching $M_{\Gamma}$ of the graph $G(Q)$ – a disjoint set of edges. For Hamiltonian cycle this matching is perfect, that is it covers all the vertices (in graph theory perfect matchings are also called $1$-factors). Then the preimage of $M_{\Gamma}$ in $N(Q,\widetilde{\Lambda}_{\Gamma})$ and in $S^{3}$ is a disjoint set of circles $C_{\Gamma}$, each circle glued of two copies of the corresponding edge. This link is the branch set of the covering. The detailed description of this link is given in [G24]. In [VM99S2] this construction was generalized from Hamiltonian cycles to Hamiltonian theta-graphs and Hamiltonian $K_{4}$-subgraphs. We will discuss the corresponding links in Section 7.
@@ 3. Main construction
Let $P$ be an ideal right-angled hyperbolic $3$-polytope. Each nonselfcrossing Eulerian cycle $\gamma$ corresponds to a link $C_{\gamma}\subset S^{3}$ consisting of $\{\text{vertices of }P\}$ circles such that its complement $S^{3}\setminus C_{\gamma}$ is homeomorphic to the hyperbolic manifold $N(P)$ glued of $4$ copies of $P$. Moreover, $C_{\gamma}$ is the branch set of the $2$-fold branched covering $N(Q(P,\gamma),\widetilde{\Lambda}_{\Gamma_{\gamma}})\to S^{3}$.
@@
Remark 3.5.
The transition from a Hamiltonian cycle $\Gamma$ on a cubic (that is $3$-valent) graph $G$ to the nonselfcrossing Eulerian cycle in the graph $G/M_{\Gamma}$ obtained by contracting the edges of the perfect matching $M_{\Gamma}$ was used in [BP17] to give a new characterisation of cubic Hamiltonian graphs having a perfect matching.
@@
Example 3.6 can be generalized as follows. It is known that any antiprism $A(k)$ ($A(3)$ is the octahedron) is an ideal right-angled $3$-polytope (see [V17]). In Fig 3 we show two different nonselfcrossing Eulerian cycles on this polytope. The hyperbolic structure on the complement to the link corresponding to the left cycle is exactly the structure defined by W.P.Thurston in the complement to the $(2k)$-link chain [T02, Example 6.8.7]. We learn this example due to the lectures by A.Yu.Vesnin and the diploma works by D.V. Chepakova [C23] and D.A. Tsygankov [T25]. The case of $A(3)$ is also mentioned in [V17, Section 5.1].
@@ 4. Edge-twists and the existence of nonselfcrossing Eulerian cycles
This follows from Examples 3.6, 3.7, and 3.8. ∎
@@ 5. Transformations of nonselfcrossing Eulerian cycles
Corollary 5.4.
Each circle of the link $C_{\Gamma}$ corresponding to a Hamiltonian cycle $\Gamma$ on a simple $3$-polytope $Q$ is linked to at least one other circle of $C_{\Gamma}$.
Construction 5.5 (Transformation of a nonselfcrossing Eulerian cycle along conjugated vertices).
@@ 6. Links associated to Hamiltonian cycles on right-angled $3$-polytopes
Remark 6.2.
In [BD25] the contraction of edges of right-angled hyperbolic polytopes producing right-angled polytopes of finite volume is discussed. In particular, the antiprism $A_{n}$ is obtained by a contraction of a perfect matching of the Löbell polytope ($n$-barrel) $L_{n}$. The inverse operation corresponds to the hyperbolic Dehn filling.
@@
A fullerene is a simple $3$-polytope with only pentagonal and hexagonal faces. It is known that any fullerene is a right-angled hyperbolic polytope and the dodecahedron is the fullerene with minimal number of facets (see more details in [E19]). It was shown by F. Kardoš [K20] that any fullerene has a Hamiltonian cycle. Each Hamiltonian cycle on a fullerene corresponds to a hyperbolic link with hyperbolic $2$-fold branched covering space.
@@
Now consider right-angled hyperbolic $3$-polytopes of finite volume. Each finite vertex of such a polytope has valency $3$, and ideal vertices have valency $4$. Using Andreev’s theorem it can be shown that cutting off ideal vertices defines a bijection between combinatorial types of right-angled hyperbolic $3$-polytopes of finite volume and almost Pogorelov polytopes different from the $4$-prism (cube) and the $5$-prism (see [DO01, Theorem 10.3.1] and [E19, Theorem 6.5]). Moreover, all quadrangles of the resulting polytope arise from ideal vertices. A simple $3$-polytope is called an almost Pogorelov polytope, if it is different from the simplex, has no $3$-belts, and any its $4$-belt surrounds a quadrangular facet. The simple polytope with $8$ facets drawn in the center at the bottom in Fig. 2 we will denote $P_{8}$. It has a nontrivial $4$-belt consisting of pentagons and surrounding two quadrangles on each side. It was shown in [E19] that this polytope has some properties similar to properties of almost Pogorelov polytopes.
@@
Let $\Gamma$ be a Hamiltonian cycle on an almost Pogorelov $3$-polytope $Q$ or the polytope $P_{8}$.
  1. (1)
    Then shrinking to points the edges of the perfect matching $M_{\Gamma}$ complementary to $\Gamma$ gives an ideal right-angled polytope $P$ if and only if each quadrangle of $Q$ has three edges in $\Gamma$. In this case the polytope $Q$ has an induced nonselfcrossing Eulerian cycle $\gamma$ such that $\Gamma=\Gamma_{\gamma}$. For the link $C_{\gamma}$ the complement $S^{3}\setminus C_{\gamma}$ is hyperbolic.
  2. (2)
    The cube $I^{3}$ and the $5$-prism $M_{5}\times I$ do not have Hamiltonian cycles with the above condition. The polytope $P_{8}$ up to combinatorial symmetries has a unique Hamiltonian cycle with the above condition shown in Fig. 2.
  3. (3)
    For $Q\notin\{I^{3},M_{5}\times I,P_{8}\}$ the $2$-fold branched covering space corresponding to $C_{\Gamma}$ becomes hyperbolic after cutting along incompressible Klein bottles corresponding to quadrangles of $Q$. For $Q=Q_{8}$ the $2$-fold branched covering space splits into two manifolds with geometry $\mathbb{L}^{2}\times\mathbb{R}$ after cutting along the incompressible torus corresponding to the $4$-belt consisting of pentagons.
@@
For each quadrangle of $Q$ the Hamiltonian cycle contains either two its opposite edges, or three of its edges. In the first case after shrinking the quadrangle becomes a bigon and we obtain no polytope. If each quadrangle intersects $\Gamma$ by three edges, then it intersects at vertices two edges of the complementary perfect matching, and by  [E19, Corollary 12.31] cutting off all the edges of the matching gives an almost Pogorelov polytope different from the $4$- and the $5$-prism and produces all its quadrangles. For $Q=P_{8}$ to use [E19, Corollary 12.31] we additionally need to check that $\Gamma$ does not contain all the four common edges of pentagons. But in this case it can not have $3$ edges on each quadrangle, so the additional condition holds. Now shrinking quadrangles to points we obtain an ideal right-angled hyperbolic $3$-polytope $P$. The same polytope we obtain by shrinking all the edges of the matching. Thus, item (1) is proved.
Item (2) follows from the direct enumeration of Hamiltonian cycles on $I^{3},M_{5}\times I$ and $P_{8}$.
By [E22a, Theorem 4.12] for $Q\notin\{I^{3},M_{5}\times I,P_{8}\}$ its quadrangles correspond to incompressible Klein bottles in $N(Q,\widetilde{\Lambda}_{\Gamma})$ such that the complement to their union has a complete hyperbolic structure of finite volume. Also by this Theorem for $Q=P_{8}$ the $4$-belt consisting of pentagons corresponds to an incompressible torus in $N(Q,\widetilde{\Lambda}_{\Gamma})$ such that its complement consists of two manifolds with geometry $\mathbb{L}^{2}\times\mathbb{R}$. This proves item (3). ∎
@@
To characterise ideal right-angled $3$-polytopes corresponding to Hamiltonian cycles on (a) compact right-angled hyperbolic $3$-polytopes (b) right-angled hyperbolic $3$-polytopes of finite volume.
@@
In [E19, Theorem 9.17] is was proved that any ideal right-angled hyperbolic $3$-polytope $P$ can be obtained from an almost Pogorelov polytope or the polytope $P_{8}$ by a contraction of edges of a perfect matching such that no quadrangle contains two edges of the matching. Nevertheless, the complement to this matching may be not a Hamiltonian cycle, but a union of cycles containing all the vertices of the polytope.